When the sun's angles of elevation is 30degrees the shadow of the post is 6 ft longer then when the angle is 45 degrees . find the height of the post.
Can you please include the work of how the question was answered.
Thank you!

Answer by stanbon(75887)

When the sun's angle of elevation is 30 degrees the shadow of the post is 6 ft longer then when the angle is 45 degrees . find the height of the post.
---
Let the height of the post be "x" ft.
Then the shadow at 45 degrees is "x" ft.
-----
Shadow at 30 degrees will be x+6 ft.
----
You have a 30/60 degree right triangle
with legs x and x+6.
--
The hypotenuse will be 2*the side opposite the 30 degree angle.
hypotenuse = 2x
---
The hypotenuse will also be sqrt[x^2 + (x+6)^2]
=====
Solve for "x":
2x = sqrt[x^2 + (x+6)^2]
---
4x^2 = x^2 + x^2+12x+36
2x^2-12x-36 = 0
----
x^2 - 6x - 18 = 0
x = [6 +- sqrt[36-4*1*-18]/(2)
---
Positive solution:
x = [6 + 10.39]/2
---
x = 8.195 ft.
==================
Cheers,
Stan H.